P(X = x) = k/x+1; where x=1,2,3
Evaluate k.
1
Expert's answer
2020-12-21T18:00:58-0500

P(X=1)+P)X=2)+P(X=3)=1

"\frac{k}{1+1}+\frac{k}{2+1}+\frac{k}{3+1}=1"

"k(\frac{1}{2}+\frac{1}{3}+\frac{1}{4})=1"

"k\frac{13}{12}=1"

"k=\frac{12}{13}"